Fermi National Laboratory
 

Accelerator Update for Jun 13, 2006 - June 14, 2006

 

Tuesday June 13

The day shift began with Operations monitoring store 4759 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Booster, Main Injector (MI), and Recycler conducting studies, and with NuMI and MiniBooNE taking beam.

Operations began preparing the Recycler at 12:18 PM to stash antiprotons.

Earlier this morning, Operations began sending beam to Switchyard (SY). A strange reading on the SY dump beam monitor, at 1:07 PM, required Operations to halt beam and investigate. The reading was false and beam resumed at 1:15 PM.

The stash ended and Pbar resumed stacking at 1:21 PM.

Operations began preparing Pbar for an access at 1:25 PM.

Operations prepared Booster for work and the H- Source for an access at 1:27 PM. This will halt beam to NuMI and MiniBooNE.

TeV experts began an end-of-store study at 2:02 PM.

The Booster and Source work finished. NuMI and MiniBooNE resumed taking beam at 2:50 PM.

Operations began turning Pbar back on at 2:57 PM. Pbar resumed stacking at 4:20 PM. The reason for the delay between turn on and stacking was due to two problems with a power supply (A:QT).

Operations terminated store 4759 at 5:03 PM.

D0 reported at 5:15 PM that they were powering down their solenoid for about an hour to calibrate.

At 5:33 PM, Operations reported that Booster beam position problems were holding off the TeV shot. Booster experts investigated and found a problem with the paraphase curve.

Operations put the TeV into shot setup around 6:16 PM.

Operations established store 4760 at 8:28 PM with an initial luminosity of 17.0E30. TeV experts found a separator at the wrong polarity, which explained the low luminosity reading. Experts changed the polarity, re-scraped the beam, and came up with a luminosity of around 51E30.

D0’s calibration ended and they began powering up their magnets at 10:05 PM.

Wednesday June 14

The midnight shift began with Operations monitoring store 4760 and a small stash, with Pbar stacking, with CDF and D0 using the colliding beam, with MI, Recycler, and SY conducting studies, and with MiniBooNE and NuMI taking beam.

At 12:04 AM, the NuMI Horn tripped off. It reset okay.

The Recycler began preparing to stash at 3 AM. The stash ended and Pbar resumed stacking at 3:56 AM.

At 5:37 AM, CDF reported that their silicon was off. They will leave it off until after their 8 AM meeting.

At 6:05 AM, Operations reported that the NuMI beam permit had tripped off due to Lambertson low level limit (I:LAM61). Operations called in an expert.

The Plans for Wednesday

The plans for today are to keep store 4760 in until this evening. The Run Coordinator would like many transfers to the Recycler so that there is a big stash for the next TeV shot setup. He also wants the antiprotons and proton intensities to increase.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/14/2006    email Fermilab